OpenStudy (anonymous):

A manufacturer finds that the revenue generated by selling x units of a certain commodity is given by the function R(x) = 100x − 0.2x2, where the revenue R(x) is measured in dollars. What is the maximum revenue, and how many units should be manufactured to obtain this maximum?

OpenStudy (anonymous):

The extrema of a function happen when the derivative is equal to zero. So first we calculate the derivative: \[R \prime(x) = 100 − 0.4x\] And then we solve: \[0 = 100 − 0.4x\] \[0.4x = 100\] x=250 units Now we can substitute back into the original equation to calculate the revenue: \[R(250) = 100(250) − 0.2(250^{2})\] R(250)=$12500